24. A material is described as ‘not easy to scratch or indent’. The material is best described as
A hard B plastic C stiff D tough
25. A spring extends by 10 cm when a force of 8.0 N is applied. The limit of proportionality is not exceeded. Two of these springs are arranged side by side and a force of 4.0 N is applied. The extension for this arrangement of springs is
A 2.5 cm
B 5.0 cm
C 10 cm
D D 20 cm
1
Expert's answer
2020-10-02T07:15:05-0400

24. Hardness is a measure of how easily a material can be scratched or indented. Thus, the answer is A


25. If the spring extends by 10 cm = 0.1m when a force of 8.0 N is applied, then its spring constant is:


"k = \dfrac{F}{x} = \dfrac{8N}{0.1m} = 80\dfrac{N}{m}"

When springs are attached in parallel, their spring constans add up. Thus, the extension for this arrangement of springs is:


"x = \dfrac{F}{2k} = \dfrac{4N}{2\cdot 80\dfrac{N}{m}} = 0.025m = 2.5 cm"

Thus, the answer is A.


Answer. 24) A, 25) A.